# Designing an Optimal Sensor Network via Minimizing Information Loss

### v1

@@ 1 Introduction
Problems in optimal experimental design have been classically studied in frequentist and Bayesian paradigms, resulting in a mature theory with hundreds of practical applications. However, these approaches typically rely on access to the true data or make model-free decisions. In parallel to the statistical development of experimental design, the proliferation of vast computational resources has profoundly impacted statistics and the sciences alike. For Bayesian statistics in particular, computational techniques such as Markov chain Monte Carlo and variational inference have led to a more widespread adoption of the Bayesian viewpoint within the broader statistics community (martin2020computing). Increased computational resources have also spurred the burgeoning field of Bayesian deep learning (papamarkou2024position). In the sciences, complex, large-domain, physics-based numerical simulations represent one innovation unlocked by computational power – for instance, in the climate sciences with the development of coupled earth system models (golaz2019doe) and global cloud-resolving models (satoh2019global), and in protein folding research with the development of simulation models of all-atom molecular dynamics (lane2013milliseconds).
@@
The interplay between these two types of simulations — statistical and physics-based — is important in practice, as their capabilities are often complementary. Asymmetry in capabilities is often attributable to interpolation or extrapolation from data observed by sensors; whereas statistical techniques excel at interpolating observations, physical simulations are much better suited to extrapolating outside of a limited observation dataset, as their predictions are carefully constrained by scientific knowledge. [^1] We note that while statistical models and physical simulations are not mutually exclusive, incorporating probabilistic estimation in physical simulations may be scientifically and/or computationally infeasible. Thus, for our purposes, we treat them as effectively separate objects. Returning to the climate sciences, statistical approaches are well suited to smoothing spatiotemporal observations (pinder2023developments), but generic approaches tend to produce high degrees of uncertainty at timescales larger than several hours (lin2005precipitation).
@@
In this work, we use data from physics-based models to determine the optimal locations for observing a spatiotemporal process using sensors. We assume that the physics-based models in question are, in some way, “accurate enough” for this task (an assumption that we formalize later), but are, in other ways, “too inaccurate,” thus driving a need to evaluate them against observations taken by sensors. To find the optimal locations for sensors, we will utilize recent advances in the variational inference of spatiotemporal Gaussian processes (GPs), which allow for efficient variational learning when spatial inducing points are fixed with respect to time and the spatio-temporal kernel is separable (i.e., decomposes as a product of a temporal and spatial kernel). The use of variational inference allows us to motivate the proposed method from an information-theoretic perspective in terms of “information loss,” while the use of physics-based simulation data can be interpreted as incorporating an “inductive bias” into our model.
@@
This inductive bias is particularly relevant, as any model-based approach to optimal experimental design necessarily relies on the quality of the model. Using physics-based simulations and variational inference allows us to create statistical models that maximally reconstruct posterior distributions under the best current knowledge available, rather than a more ignorant prior.
@@
At this point, it is prudent to contextualize our method within the more general framework of Bayesian experimental design (BED). More precisely, BED deals with finding an “optimal experimental design,” which may include selecting the locations where a spatiotemporal process is observed, guided by information-theoretic criteria (chaloner1995Bayesian). However, BED typically targets intractable quantities, which makes optimization extremely difficult. In fact, some of the most popular utilities used in BED (e.g., the expected information gain) are doubly intractable, meaning they involve nested expectations over intractable posteriors. Therefore, even standard Monte Carlo methods cannot be applied in general (rainforth2024modern).
@@
As a result, much of the effort in BED is related to computational approaches, such as improved nested Monte Carlo estimators (rainforth2018nesting; iqbal2024nesting) or approximate inference schemes (foster2019variational). An extremely common approximation is to optimize in a greedy, sequential manner, in which case BED is known as Bayesian adaptive design (BAD). However, in the case of selecting locations to observe spatiotemporal processes, this greedy selection may lead to undesirable behaviors.
@@
This paper is structured as follows: in Section 2, we provide background on BED using surrogate models and particularly motivate our notion of optimality. In Section 3, we introduce our “minimum information loss (MIL)” optimality criterion and discuss the properties of our method. In Section 4, we present a brief review of Gauss-Markov regression, focusing in particular on the recently proposed spatiotemporal variational Gaussian process (STVGP) (hamelijnck2021spatio), and show how it enables efficient computation of MIL-optimal solutions in spatiotemporal settings. We discuss more intimate connections with “modern” OED methods, including other variational ones. Finally, we show the performance of our method on a practical case study in air temperature modeling in Section 6, and conclude with some discussion and directions for future work in Section 7. We also derive several novel methods for continuous sensor placement as baselines using STVGPs, explicated in Appendix C.
@@ 2 Model-Based Experimental Design
In this section, we overview some existing methods for optimal experimental design — with a particular emphasis on Bayesian techniques — and use this to motivate our proposed method. For simplicity, we first focus on the case where some real, scalar process $Y(\boldsymbol{\mathbf{x}})\in\mathbb{R}$ is observed, indexed by some quantity $\mathbf{x}$ that belongs to a compact subset $\mathcal{X}\subset\mathbb{R}^{d_{X}}$. We outline existing methods related to the expected information gain (Section 2.1), followed by a more general utility-based experimental design (Section 2.3). An expert in Bayesian experimental design may consider skipping to Section 3 after familiarizing themselves with our notation.
@@ 2.1 Existing Approaches in Bayesian Experimental Design
In general, Bayesian experimental design aims to find the design $\xi\in\Xi$ that allows us to collect data $D\in\mathcal{D}$ and reduce our uncertainty about the parameter of interest $\theta\in\Theta$, through the use of a utility function $U(\xi,D,\theta)$ and a posterior distribution $p(\theta|D,\xi).$ [^2] Since these methods derive their optimality through a statistical model, we will call them “model-based” (juxtaposed to model-free approaches such as Latin Hypercube sampling). For learning about the process $Y({\bf x})$, the design problem refers to the task of selecting a suitable set of $n$ inputs $\boldsymbol{\mathbf{X}}_{n}\in\mathcal{X}^{n}$ where a process ${Y}(\cdot)$ is observed, resulting in observations $\boldsymbol{\mathbf{Y}}_{n}$, such that these observations are as informative as possible about the process, especially regarding unobserved outputs $\boldsymbol{\mathbf{Y}}_{\text{test}}\triangleq\mathbf{Y}(\mathbf{X}_{\text{test}})$. In frequentist analysis, the notion of “information” is typically related to Fisher information (pukelsheim2006optimal; box1982choice), while Bayesian approaches center on Shannon’s information theory instead (chaloner1995Bayesian; rainforth2024modern).
In a Bayesian context, the most commonly used criterion for optimality is differential entropy or (negative) information. Recall the differential cross-entropy of a random variable $A$ with pdf $p_{A}(x)$ and another random variable $B$ with pdf $p_{B}(x)$. [^3] We will generally assume that $A$ and $B$ have densities with respect to the Lebesgue measure and are mutually dominating measures. It is given by
@@
$\operatorname{EIG}(\mathbf{X}_{n})\triangleq H(\boldsymbol{\mathbf{Y}}_{\text{test}})-H(\boldsymbol{\mathbf{Y}}_{\text{test}}\,|\,\mathbf{Y}_{n},\mathbf{X}_{n}).$ (3)
@@
is a nested expectation with respect to the generally intractable marginal predictive distribution $p(\boldsymbol{\mathbf{Y}}_{\text{test}},{\bf Y}_{n}\,|\,\mathbf{X}_{n})$. It is thus a “doubly-intractable quantity,” which cannot be estimated with simple Monte Carlo methods (rainforth2024modern).
@@ 2.2 Bayesian Adaptive Design
To address the computational intractability of evaluating EIG, one often resorts to greedy optimization or simple statistical models. The greedy optimization approach, known as Bayesian adaptive design (BAD), maximizes the EIG at index $n$ using previously selected sensor locations $\mathbf{X}_{n-1}$. This process identifies a new sensor location $\mathbf{x}_{n}$ and forms the new set of sensor locations $\mathbf{X}_{n}=\mathbf{X}_{n-1}\cup\{\mathbf{x}_{n}\}$, where location $\mathbf{x}_{n}$ is the location that maximizes $\operatorname{EIG}(\mathbf{X}_{n})$.
@@
If $\mathbf{X}_{n}$ and $\mathbf{X}_{\text{test}}$ are chosen to form a partition of $\mathcal{X}$, i.e., $\mathcal{X}=\mathbf{X}_{n}\sqcup\mathbf{X}_{\text{test}}$, then Eq. 5 becomes the mutual information $I(\mathbf{Y}_{\mathcal{X}\setminus\mathbf{X}_{n}};\boldsymbol{\mathbf{Y}}_{n})$. While discrete optimization of this objective is still intractable, krause2008near show that greedy optimization of the mutual information provides near-optimal sampling locations, with (11/e)(1-1/e) convergence. However, the greedy algorithm is not guaranteed to converge if the choice in $\mathbf{X}_{\text{test}}$ is independent of the choice in $\mathbf{X}_{n}$, i.e., if the set of test locations is fixed and does not depend on the sensor network.
@@
The challenge in providing a tractable algorithm in krause2008near is that the optimization problem is discrete, leading to combinatorial complexity. To select $k$ sensor locations out of $n$ candidates, the greedy algorithm under a full GP model incurs a total computational cost of $\mathcal{O}(kn^{4})$. This cost arises from two sources: at each of the $k$ greedy selection steps, the algorithm must evaluate the utility (mutual information) for up to $n$ remaining candidates, which results in $\mathcal{O}(kn)$ utility evaluations. Each utility computation, in turn, generally requires $\mathcal{O}(n^{3})$ time due to matrix inversions or log-determinant computations associated with GP inference. Thus, the total runtime scales as $\mathcal{O}(kn^{4})$ when performed naively with full GP models. Even in the most generous case, their algorithm must explore $\mathcal{O}(kn)$ combinations of locations, which may still prove prohibitively expensive. On the contrary, our proposed approach is continuous in nature and therefore does not require a discretization of $\mathcal{X}$ for optimization.
@@ 2.3 Utility-Based Optimal Experimental Design
Taking a more global view, model-based design criteria may maximize the expected value of some utility other than the information gain. Many can be interpreted under the umbrella of Bayes-optimal design, where the optimal observations maximize a utility function $U(\cdot)$ (chaloner1995Bayesian),
@@
For example, we may rewrite the EIG in terms of the KL divergence between the posterior and prior of $\mathbf{Y}_{\text{test}}$ (cover1999elements),
$\operatorname{EIG}({\bf X}_{n})=\mathbb{E}_{\mathbf{Y}_{n}}\mathopen{}\mathclose{{\left[\operatorname{KL}(p(\mathbf{Y}_{\text{test}}\,|\,\mathbf{Y}_{n},\mathbf{X}_{n})\,||\,p(\mathbf{Y}_{\text{test}}))}}\right].$ (7)
@@
In fact, this was the original motivation for EIG, pioneered by lindley1956measure.
There are many other utility-based approaches that may be based on conditional entropy, posterior predictive variance, and more. Several of these criteria amount to placing different functional norms on the variance; for example, the integrated mean-square error criterion minimizes the posterior predictive variance integrated over $\mathcal{X}$, and the maximum mean-square error (MMSE) criterion minimizes the maximum posterior predictive variance (sacks1989). In either case, finding all the optimal sensor locations ($\mathbf{X}^{\text{opt}}_{n}$) in “one shot” is generally difficult, and greedy sequential approaches are used instead (gramacy2020surrogates, Ch. 6). We discuss some additional utility-based Bayesian OED methods, most notably maximum entropy sampling (MES) in Appendix A.
@@ 3.1 Motivating and Defining the MIL Criterion
The model-based design criteria discussed in Section 2.1 and their sequential implementations have several drawbacks that we attempt to address. One notable drawback of the aforementioned design criteria (and MES, in particular) is that they tend to resemble naïve space-filling designs that do not depend on the data or statistical models (gramacy2020surrogates, Ch. 4). Intuitively, this can be undesirable if the observed process $Y(\cdot)$ is spatially nonstationary, with some areas being more difficult to predict than others. Indeed, in their work with GPs, krause2008near show empirically that using nonstationary kernels while maximizing mutual information significantly improves performance and argue theoretically that stationary kernels result in space-filling designs.
@@ 3.2 Connection to Sparse Variational Gaussian Processes
To motivate this connection, we first recall the setting of sparse GP regression. Consider the scenario where we would like to fit a GP $f(\cdot)$ on a dataset $\mathcal{D}=\{(\mathbf{x}_{n},y_{n})\}_{n=1}^{N}$, where $N$ is very large. We will use $\mathbf{X}_{N}$ to refer to the inputs of $\mathcal{D}$, and $\mathbf{Y}_{N}$ to refer to the outputs. In this case, exact GP interpolation may be prohibitively expensive, as the computational complexity of posterior predictions is $\mathcal{O}(N^{3})$ (rasmussen_and_williams). To simplify computations, we introduce a set of $M\ll N$ inducing points $\mathbf{Z}_{M}=\{{\bf z}_{m}\}_{m=1}^{M}\subset\mathcal{X}$. Denoting the function values at $\mathbf{X}_{N}$ as $\mathbf{f}_{N}$ and the function values at $\mathbf{Z}_{M}$ as $\mathbf{u}_{M}$, the posterior predictive of the full GP is then approximated as
@@
We observe a superficial resemblance between optimizing the objective of Section 3.2 and the various model-based utilities discussed in Section 2.1: Section 3.2 consists of a regularization term based on the prior covariance matrix and a model fit term that encourages agreement between $\mathbf{Y}_{N}$ and the prior predictive mean. Nevertheless, we focus more specifically on the sparse variational approach (titsias2009variational), which is known to guard against overfitting and has become the standard framework for sparse GPs.
@@
which is closest to the true posterior (13) in terms of the KL divergence; hence, it being obtained via
@@
In the case of Gaussian likelihoods with variance $\sigma_{\text{obs}}^{2}$, the minimization problem is equivalent to the maximization of the negative evidence lower bound (ELBO),
[…]
$\overset{+}{=}-{\bf Y}_{N}\mathopen{}\mathclose{{\left(\mathbf{Q}_{NN}+\sigma^{2}_{\text{obs}}\mathbf{I}}}\right)^{-1}{\bf Y}_{N}-\operatorname{log\,det}\mathbf{Q}_{NN}-\frac{1}{2\sigma_{\text{obs}}^{2}}\operatorname{Tr}(\mathbf{K}_{\text{post}}({\bf X}_{N})),$ (17)
@@
where $\mathbf{K}_{NN}$, $\mathbf{K}_{MM}$ are kernel matrices computed using ${\bf X}_{N}$ and ${\bf Z}_{M}$, respectively, and $\mathbf{K}_{MN}=\mathbf{K}_{NM}^{\top}$ is the rectangular matrix containing cross-covariances. Note that Eq. 17 is only a function of ${\bf Z}_{M}$ since, in the case of a Gaussian variational family, the optimal expression for the variational distribution is available in closed form, $q_{\texttt{opt}}({\bf u}_{M}|{\bf Y}_{N},{\bf X}_{N},{\bf Z}_{M})=\mathcal{N}\mathopen{}\mathclose{{\left({\bf u}_{M}|\boldsymbol{\mu},\mathbf{A}}}\right)$, for known $\boldsymbol{\mu}$ and $\mathbf{A}$.
@@
As noted in our derivations, the use of MIL-optimality requires a simulator that is somehow “accurate enough” for sensor placement design, but “not accurate enough” for full-scale scientific policy. We argue here that this situation is not uncommon in practice and formalize this notion for MIL-optimality.

### v2

@@ 1 Introduction
Problems in optimal experimental design (OED) have been classically studied in frequentist and Bayesian paradigms, resulting in a mature theory with hundreds of practical applications. However, these approaches typically rely on access to the true data or make model-free decisions. In parallel to the statistical development of experimental design, the proliferation of vast computational resources has profoundly impacted statistics and the sciences alike. For Bayesian statistics in particular, computational techniques such as Markov chain Monte Carlo and variational inference have led to a more widespread adoption of the Bayesian viewpoint within the broader statistics community (Martin et al., 2020). Increased computational resources have also spurred the burgeoning field of Bayesian deep learning (Papamarkou et al., 2024). In the sciences, complex, large-domain, physics-based numerical simulations represent one innovation unlocked by computational power – for instance, in the climate sciences with the development of coupled earth system models (Golaz et al., 2019) and global cloud-resolving models (Satoh et al., 2019), and in protein folding research with the development of simulation models of all-atom molecular dynamics (Lane et al., 2013).
@@
The interplay between these two types of simulations — statistical and physics-based — is important in practice, as their capabilities are often complementary. Asymmetry in capabilities is often attributable to interpolation or extrapolation from data observed by sensors; whereas statistical techniques excel at interpolating observations, physical simulations are much better suited to extrapolating outside of a limited observation dataset, as their predictions are carefully constrained by scientific knowledge. [^1] We note that while statistical models and physical simulations are not mutually exclusive, incorporating probabilistic estimation in physical simulations may be scientifically and/or computationally infeasible. Thus, for our purposes, we treat them as effectively separate objects. Returning to the climate sciences, statistical approaches are well-suited to smoothing spatiotemporal observations, but generic approaches tend to produce high degrees of uncertainty at longer forecasting timescales (Pinder, 2023).
@@
In this work, we use data from physics-based models to determine the optimal locations for observing a spatiotemporal process using sensors. We assume that the physics-based models in question are, in some way, “accurate enough” for this task (an assumption that we formalize later), but are, in other ways, “too inaccurate,” thus driving a need to evaluate them against observations taken by sensors. To find the optimal locations for sensors, we will utilize recent advances in the variational inference of spatiotemporal Gaussian processes (GPs), which allow for efficient variational learning when spatial inducing points are fixed with respect to time and the spatiotemporal kernel is separable (i.e., decomposes as a product of a temporal and spatial kernel). The use of variational inference allows us to motivate the proposed method from an information-theoretic perspective in terms of “information loss,” while the use of physics-based simulation data can be interpreted as incorporating an “inductive bias” into our model.
@@
This inductive bias is particularly relevant, as any model-based approach to OED necessarily relies on the quality of the model. Using physics-based simulations and variational inference allows us to create statistical models that maximally reconstruct posterior distributions under the best current knowledge available, rather than a more ignorant prior.
@@
At this point, it is prudent to contextualize our method within the more general framework of Bayesian experimental design (BED). More precisely, BED deals with finding an “optimal experimental design,” which may include selecting the locations where a spatiotemporal process is observed, guided by information-theoretic criteria (Chaloner and Verdinelli, 1995). However, BED typically targets intractable quantities, which makes optimization extremely difficult. In fact, some of the most popular utilities used in BED (e.g., the expected information gain) are doubly intractable, meaning they involve nested expectations over intractable posteriors. Therefore, even standard Monte Carlo methods cannot be applied in general (Rainforth et al., 2024).
@@
As a result, much of the effort in BED is related to computational approaches, such as improved nested Monte Carlo estimators (Rainforth et al., 2018) or approximate, amortized inference schemes (Foster et al., 2019). An extremely common approximation is to optimize in a greedy, sequential manner, in which case BED is known as Bayesian adaptive design (BAD). However, in the case of selecting locations to observe spatiotemporal processes, this greedy selection may lead to undesirable behaviors.
@@
This paper is structured as follows: in Section 2, we provide background on BED using surrogate models and particularly motivate our notion of optimality. In Section 3, we introduce our “minimum information loss (MIL)” optimality criterion and discuss the properties of our method. In Section 4, we present a brief review of Gauss-Markov regression, focusing in particular on the recently proposed spatiotemporal variational Gaussian process (STVGP) (Hamelijnck et al., 2021), and show how it enables efficient computation of MIL-optimal solutions in spatiotemporal settings. We discuss more intimate connections with “modern” OED methods, including other variational ones. Finally, we show the performance of our method on a practical case study in air temperature modeling in Section 6, and conclude with a discussion and directions for future work in Section 7. We also derive several novel methods for continuous sensor placement as baselines using STVGPs, as explicated in Appendix C (Waxman et al., 2025).
@@ 2 Model-Based Experimental Design
In this section, we overview some existing methods for OED — with a particular emphasis on Bayesian techniques — and use this to motivate our proposed method. For simplicity, we first focus on the case where some real, scalar process $Y(\boldsymbol{\mathbf{x}})\in\mathbb{R}$ is observed, indexed by some quantity $\mathbf{x}$ that belongs to a compact subset $\mathcal{X}\subset\mathbb{R}^{d_{X}}$. We outline existing methods related to the expected information gain (Section 2.1), followed by a more general utility-based experimental design (Section 2.3). An expert in Bayesian experimental design may consider skipping to Section 3 after familiarizing themselves with our notation.
@@ 2.1 Existing Approaches in Bayesian Experimental Design
In general, Bayesian experimental design aims to find the design $\xi\in\Xi$ that allows us to collect data $D\in\mathcal{D}$ and reduce our uncertainty about the parameter of interest $\theta\in\Theta$, through the use of a utility function $U(\xi,D,\theta)$ and a posterior distribution $p(\theta|D,\xi).$ [^2] Since these methods derive their optimality through a statistical model, we will call them “model-based” (juxtaposed to model-free approaches such as Latin Hypercube sampling). For learning about the process $Y({\bf x})$, the design problem refers to the task of selecting a suitable set of $n$ inputs $\boldsymbol{\mathbf{X}}_{n}\in\mathcal{X}^{n}$ where a process ${Y}(\cdot)$ is observed, resulting in observations $\boldsymbol{\mathbf{Y}}_{n}$, such that these observations are as informative as possible about the process, especially regarding unobserved outputs $\boldsymbol{\mathbf{Y}}_{\text{test}}\triangleq\mathbf{Y}(\mathbf{X}_{\text{test}})$. In frequentist analysis, the notion of “information” is typically related to Fisher information (Pukelsheim, 2006; Box, 1982), while Bayesian approaches center on Shannon’s information theory instead (Chaloner and Verdinelli, 1995; Rainforth et al., 2024).
In a Bayesian context, the most common criterion for optimality is differential entropy or (negative) information. Recall the differential cross-entropy of a random variable $A$ with pdf $p_{A}(x)$ and another random variable $B$ with pdf $p_{B}(x)$. [^3] We will generally assume that $A$ and $B$ have densities with respect to the Lebesgue measure and are mutually dominating measures. is given by
@@
$\operatorname{EIG}(\mathbf{X}_{n})\triangleq\mathbb{E}_{\boldsymbol{\mathbf{Y}}_{n}\,|\,\mathbf{X}_{n}}\mathopen{}\mathclose{{\left[H(\boldsymbol{\mathbf{Y}}_{\text{test}})-H(\boldsymbol{\mathbf{Y}}_{\text{test}}\,|\,\mathbf{Y}_{n},\mathbf{X}_{n})}}\right].$ (3)
@@
is a nested expectation with respect to the generally intractable marginal predictive distribution $p(\boldsymbol{\mathbf{Y}}_{\text{test}},{\bf Y}_{n}\,|\,\mathbf{X}_{n})$. It is thus a “doubly-intractable quantity,” which cannot be estimated with simple Monte Carlo methods (Rainforth et al., 2024).
@@ 2.2 Bayesian Adaptive Design
To address the computational intractability of evaluating EIG, one often resorts to greedy optimization or simple statistical models. The greedy optimization approach, known as Bayesian adaptive design (BAD), maximizes the EIG at index $n$ using previously selected sensor locations $\mathbf{X}_{n-1}$. This process identifies a new sensor location $\mathbf{x}_{n}$ and forms a new set of sensor locations $\mathbf{X}_{n}=\mathbf{X}_{n-1}\cup\{\mathbf{x}_{n}\}$, where location $\mathbf{x}_{n}$ is the location that maximizes $\operatorname{EIG}(\mathbf{X}_{n})$.
@@
If $\mathbf{X}_{n}$ and $\mathbf{X}_{\text{test}}$ are chosen to form a partition of $\mathcal{X}$, i.e., $\mathcal{X}=\mathbf{X}_{n}\sqcup\mathbf{X}_{\text{test}}$, then Eq. 5 becomes the mutual information $I(\mathbf{Y}_{\mathcal{X}\setminus\mathbf{X}_{n}};\boldsymbol{\mathbf{Y}}_{n})$. While discrete optimization of this objective is still intractable, Krause et al. (2008) show that greedy optimization of the mutual information provides near-optimal sampling locations, with (11/e)(1-1/e) convergence. However, the greedy algorithm is not guaranteed to converge if the choice in $\mathbf{X}_{\text{test}}$ is independent of the choice in $\mathbf{X}_{n}$, i.e., if the set of test locations is fixed and does not depend on the sensor network.
@@
The challenge in providing a tractable algorithm in Krause et al. (2008) is that the optimization problem is discrete, leading to combinatorial complexity. To select $k$ sensor locations out of $n$ candidates, the greedy algorithm under a full GP model incurs a total computational cost of $\mathcal{O}(kn^{4})$. This cost arises from two sources: at each of the $k$ greedy selection steps, the algorithm must evaluate the utility (mutual information) for up to $n$ remaining candidates, which results in $\mathcal{O}(kn)$ utility evaluations. Each utility computation, in turn, generally requires $\mathcal{O}(n^{3})$ time due to matrix inversions or log-determinant computations associated with GP inference. Thus, the total runtime scales as $\mathcal{O}(kn^{4})$ when performed naively with full GP models. Even in the most generous case, their algorithm must explore $\mathcal{O}(kn)$ combinations of locations, which may still prove prohibitively expensive. On the contrary, our proposed approach is continuous in nature and therefore does not require a discretization of $\mathcal{X}$ for optimization.
@@ 2.3 Utility-Based Optimal Experimental Design
Taking a more global view, model-based design criteria may maximize the expected value of some utility other than the information gain. Many can be interpreted under the umbrella of Bayes-optimal design, where the optimal observations maximize a utility function $U(\cdot)$ (Chaloner and Verdinelli, 1995),
@@
For example, we may rewrite the EIG in terms of the KL divergence between the posterior and prior of $\mathbf{Y}_{\text{test}}$ (Cover, 1999),
$\operatorname{EIG}({\bf X}_{n})=\mathbb{E}_{\mathbf{Y}_{n}|\mathbf{X}_{n}}\mathopen{}\mathclose{{\left[\operatorname{KL}(p(\mathbf{Y}_{\text{test}}\,|\,\mathbf{Y}_{n},\mathbf{X}_{n})\,||\,p(\mathbf{Y}_{\text{test}}))}}\right].$ (7)
@@
In fact, this was the original motivation for EIG, pioneered by Lindley (1956).
There are many other utility-based approaches that may be based on conditional entropy, posterior predictive variance, and more. Several of these criteria amount to placing different functional norms on the variance; for example, the integrated mean-square error criterion minimizes the posterior predictive variance integrated over $\mathcal{X}$, and the maximum mean-square error (MMSE) criterion minimizes the maximum posterior predictive variance (Sacks et al., 1989). In either case, finding all the optimal sensor locations ($\mathbf{X}^{\text{opt}}_{n}$) in “one shot” is generally difficult, and greedy sequential approaches are used instead (Gramacy, 2020, Ch. 6). We discuss some additional utility-based Bayesian OED methods, most notably maximum entropy sampling (MES) in Appendix A (Waxman et al., 2025)
@@ 3.1 Motivating and Defining the MIL Criterion
The model-based design criteria discussed in Section 2.1 and their sequential implementations have several drawbacks that we attempt to address. One notable drawback of the aforementioned design criteria (and MES, in particular) is that they tend to resemble naïve space-filling designs that do not depend on the data or statistical models (Gramacy, 2020, Ch. 4). Intuitively, this can be undesirable if the observed process $Y(\cdot)$ is spatially nonstationary, with some areas being more difficult to predict than others. Indeed, in their work with GPs, Krause et al. (2008) show empirically that using nonstationary kernels while maximizing mutual information significantly improves performance and argue theoretically that stationary kernels result in space-filling designs.
@@ 3.2 Connection to Sparse Variational Gaussian Processes
To motivate this connection, we first recall the setting of sparse GP regression. Consider the scenario where we would like to fit a GP $f(\cdot)$ on a dataset $\mathcal{D}=\{(\mathbf{x}_{n},y_{n})\}_{n=1}^{N}$, where $N$ is very large. We will use $\mathbf{X}_{N}$ to refer to the inputs of $\mathcal{D}$, and $\mathbf{Y}_{N}$ to refer to the outputs. In this case, exact GP interpolation may be prohibitively expensive, as the computational complexity of posterior predictions is $\mathcal{O}(N^{3})$ (Rasmussen and Williams, 2006). To simplify computations, we introduce a set of $M\ll N$ inducing points $\mathbf{Z}_{M}=\{{\bf z}_{m}\}_{m=1}^{M}\subset\mathcal{X}$. Denoting the function values at $\mathbf{X}_{N}$ as $\mathbf{f}_{N}$ and the function values at $\mathbf{Z}_{M}$ as $\mathbf{u}_{M}$, the posterior predictive of the full GP is then approximated as
@@
We observe a superficial resemblance between optimizing the objective of Section 3.2 and the various model-based utilities discussed in Section 2.1: Section 3.2 consists of a regularization term based on the prior covariance matrix and a model fit term that encourages agreement between $\mathbf{Y}_{N}$ and the prior predictive mean. Nevertheless, we focus more specifically on the sparse variational approach (Titsias, 2009), which is known to guard against overfitting and has become the standard framework for sparse GPs.
@@
which is closest to the true posterior (13) in terms of the KL divergence. This variational distribution is obtained by solving
@@
In the case of Gaussian likelihoods with variance $\sigma_{\text{obs}}^{2}$, this minimization problem is equivalent to maximizing the evidence lower bound (ELBO),
$\mathcal{L}_{V}\mathopen{}\mathclose{{\left({\bf Z}_{M}}}\right)$ $=\log\mathopen{}\mathclose{{\left[\mathcal{N}\mathopen{}\mathclose{{\left({\bf Y}_{N}\mid\mathbf{0},\sigma_{\text{obs}}^{2}\mathbf{I}+\mathbf{Q}_{NN}}}\right)}}\right]-\frac{1}{2\sigma_{\text{obs}}^{2}}\operatorname{Tr}(\mathbf{K}_{\text{post}}({\bf X}_{N})),$
$\overset{+}{=}-\frac{1}{2}{\bf Y}_{N}^{\top}\mathopen{}\mathclose{{\left(\mathbf{Q}_{NN}+\sigma^{2}_{\text{obs}}\mathbf{I}}}\right)^{-1}{\bf Y}_{N}-\frac{1}{2}\operatorname{log\,det}\mathopen{}\mathclose{{\left(\mathbf{Q}_{NN}+\sigma^{2}_{\text{obs}}\mathbf{I}}}\right)$
$\hskip 11.38092pt-\frac{1}{2\sigma_{\text{obs}}^{2}}\operatorname{Tr}(\mathbf{K}_{\text{post}}({\bf X}_{N})),$ (17)
@@
where $\mathbf{K}_{NN}$, $\mathbf{K}_{MM}$ are kernel matrices computed using ${\bf X}_{N}$ and ${\bf Z}_{M}$, respectively, and $\mathbf{K}_{MN}=\mathbf{K}_{NM}^{\top}$ is the rectangular matrix containing cross-covariances. Note that $\mathcal{L}_{V}(\mathbf{Z}_{M})$ is only a function of ${\bf Z}_{M}$ since, in the case of a Gaussian variational family, the optimal expression for the variational distribution is available in closed form, $q_{\texttt{opt}}({\bf u}_{M}|{\bf Y}_{N},{\bf X}_{N},{\bf Z}_{M})=\mathcal{N}\mathopen{}\mathclose{{\left({\bf u}_{M}|\boldsymbol{\mu},\mathbf{A}}}\right)$, for known $\boldsymbol{\mu}$ and $\mathbf{A}$.
@@
As noted in our derivations, the use of MIL-optimality requires a simulator that is “accurate enough” for sensor placement design, but “not accurate enough” for full-scale scientific policy. We argue here that this situation is not uncommon in practice and formalize this notion for MIL-optimality.